A point charge has a charge of 9.00 x 10^-11C. At what distance from the point charge is the electrical potential (a) 13.0 V? (b) 25.0 V?


1
Expert's answer
2022-03-30T13:51:09-0400

We know that

Potential

"V=\frac{kq}{r}\\r=\frac{kq}{V}\\r=\frac{9\times10^9\times9\times10^{-11}}{13}=0.0623m"

Part (b)

"V=\frac{kq}{r}\\r=\frac{kq}{V}\\r=\frac{9\times10^9\times9\times10^{-11}}{25}=0.0324m"